OpenStudy (anonymous):

evaluate the expression without using a calculator. (cos 30)^2 - (sin 30)^2

OpenStudy (anonymous):

is 30 in radians or degrees

OpenStudy (anonymous):

degress

OpenStudy (anonymous):

degrees

OpenStudy (anonymous):

use double angle formulas\[\cos (30)^2-\sin(30)^2=\cos(2*30)=\cos(60)=0.5\]

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

yes